OpenStudy (anonymous):

need help on part 2 review!!!!

OpenStudy (anonymous):

you can't assume everyone knows what you are talking about.

OpenStudy (anonymous):

Go ahead and post your question and I will try to explain.

OpenStudy (anonymous):

i know iwas about to upload da file thwen my CPU froze

OpenStudy (anonymous):

Okay. All of it or do you have something specific?

OpenStudy (anonymous):

all of it please

OpenStudy (anonymous):

Okay. Well the first problem with the table. For that one, it is simply plug and chug. So since it gives you a bunch of x and y values. You can take your answer choice and plug in the x values and see if it gives you the corresponding y values on the table.

OpenStudy (anonymous):

For the next you, when you set the equations equal you want to know where they intersect. So if it asks for the x value, simply find the x that falls directly UNDER the intersection. In this case: 2

OpenStudy (anonymous):

For the next one, the range is the y values. So, simple list the y values of the points: {2,4,5,6} 5 appears twice but you only need to list it once.

OpenStudy (anonymous):

soc on number 3 is wrong right

OpenStudy (anonymous):

No, c is correct.

OpenStudy (anonymous):

To test for symmetry about the origin. You need to see if for a point (a,b) there exists another point (-a,-b) So for this problem, plug in 1 for example, then -1. You get: (-1,2) but when you plug in positive 1 you get. (1,-2). So it is symmetric about the origin.

OpenStudy (anonymous):

ok i understand that and did i get number one correct???

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

okok and on #2 x = -2

OpenStudy (anonymous):

No, on this one x=2. Because if you go to the x directly under the intersection of the two lines (not the axes!) x is +2.

OpenStudy (anonymous):

ok thank u so much u wanna help me with the other page??

OpenStudy (anonymous):

Sure, as long as you click good answer for me :P haha

OpenStudy (anonymous):

You can just upload the page if thats easiest :)

OpenStudy (anonymous):

cool i can do dat ur a big help i have to get dis packet done or else and its up loading now

OpenStudy (anonymous):

alright, no problem at all :)

OpenStudy (anonymous):

oki think this is it

OpenStudy (anonymous):

thats the same page

OpenStudy (anonymous):

no wrong one hold on

OpenStudy (anonymous):

k here we go

OpenStudy (anonymous):

Okay. For the first one, I'm just going to go down the answer choices: For for A: Its zero is 3. Zeros are where the graph crosses the x-axis. (also known as x-intercepts). But, according to this graph you can see that it does NOT cross at 3. This is incorrect. For B: It's minimum value is -6. Here, you want to see where the lowest Y value is. And if you look all the way at the bottom, the lowest point IS -6. For C: On (1,6] the function is decreasing (i.e., negative slope). For D: The domain is all the x values for the graph. As you can see, x ranges from -4 to 6. So the WRONG one (which is NOT correct) would be answer choice A. Since zeros refers to the x intercepts NOT the y intercepts.

OpenStudy (anonymous):

For the second one, you notice that f(x) would be 3x? Take any x value and multiply it by 3 and you get the corresponding y value. So, it is ONLY positive integers (and zero). So this makes the answer choice A because it contains ONLY integers greater than or equal to zero.

OpenStudy (anonymous):

For the 3rd one, you are correct. The rate is NOT constant. (Tell me if you need an explanation)

OpenStudy (anonymous):

Then for the last one, a geometric sequence is something in the form. ar^n So in this case you have: 10,40,160,640. Factor out a 10: 10(1,4,16,64...) And you see you have 10(4^n) Which is what is underlined.

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yea gima sec im workin it out

OpenStudy (anonymous):

Okay :P

OpenStudy (anonymous):

ok ok thank u ready for the next?? ive got 7 pages to do want 2 continu??

OpenStudy (anonymous):

Haha, I'll do one more then I have to go :P

OpenStudy (anonymous):

cool ur a big help im uploadin it now

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

here we r

OpenStudy (anonymous):

Alright. For the first one, you can factor it or use the quadratic formula. I notice that -10 factors to -5 and +2. So: (x-5)(x+2). That means that your solutions are x=5 and x=-2 because (5-5)=0 and (-2+2)=0. You just have your signs switched.

OpenStudy (anonymous):

For the next one, the solutions are wherever Y is zero. So according to the chart that is x=5 and x=7. Same as the previous problem, you're setting y=0.

OpenStudy (anonymous):

For the next one. Start my multiplying both sides. So: \[(4+k)(6)=(k-2)(8)\]. Distribute. \[24+6k=8k-16\]. Add the 16 and subtract the 6k: \[40=2k\]. Divide by 2. k=20. Which you have.

OpenStudy (anonymous):

For the next one, square both sides: \[(\sqrt{3z+1})^2=5^2\]. \[3z+1=25\]. Subtract the 1 3z=24 Divide by 3. z=8. Which you also have.

OpenStudy (anonymous):

And for the last one, multiply out the left side: \[(x+3)(x+3)=x^2+6x+9=2\]. Subtract the 2. \[x^2+6x+7=0\]. This time, you have to use the quadratic formula. \[\frac{-6 \pm \sqrt{6^2-4(1)(7)}}{2(1)}\]. This gives:\[-3 \pm \frac{\sqrt{8}}{2}\]. But \[\sqrt{8}=2\sqrt{2}\]. \[-3 \pm \frac{2\sqrt{2}}{2}=-3 \pm \sqrt{2}\]. Which you have.

OpenStudy (anonymous):

ok thanks HI FIVE ur awesome man!!!!! thank u soo much i feel retricethought i might pass math now ur awesome!!!! dontlet anyone tell u otherwise

OpenStudy (anonymous):

Haha, thank you :P I'm in my senior year of college as a math/physics major :P Sorry for the lengthy explanations. I hate just giving away answers I prefer understanding :P Good luck man! I'm sure you'll do fine!

OpenStudy (anonymous):

thanks gud luck in college u will pass it for me i gotta make it 2 10th grade